\begin{thebibliography}{34}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bain et~al.(2021)Bain, Nagrani, Varol, and Zisserman]{bain2021frozen}
Max Bain, Arsha Nagrani, G{\"u}l Varol, and Andrew Zisserman.
\newblock Frozen in time: A joint video and image encoder for end-to-end
  retrieval.
\newblock In \emph{Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pp.\  1728--1738, 2021.

\bibitem[Chen et~al.(2019)Chen, D'Arcy, Liu, Fernandez, and
  Downey]{chen2019codah}
Michael Chen, Mike D'Arcy, Alisa Liu, Jared Fernandez, and Doug Downey.
\newblock Codah: An adversarially authored question-answer dataset for common
  sense.
\newblock \emph{arXiv preprint arXiv:1904.04365}, 2019.

\bibitem[Chen et~al.()Chen, Li, Yu, El~Kholy, Ahmed, Gan, Cheng, and
  Liu]{chenuniter}
Yen-Chun Chen, Linjie Li, Licheng Yu, Ahmed El~Kholy, Faisal Ahmed, Zhe Gan,
  Yu~Cheng, and Jingjing Liu.
\newblock Uniter: Universal image-text representation learning.

\bibitem[Dou et~al.(2022)Dou, Xu, Gan, Wang, Wang, Wang, Zhu, Zhang, Yuan,
  Peng, et~al.]{dou2022empirical}
Zi-Yi Dou, Yichong Xu, Zhe Gan, Jianfeng Wang, Shuohang Wang, Lijuan Wang,
  Chenguang Zhu, Pengchuan Zhang, Lu~Yuan, Nanyun Peng, et~al.
\newblock An empirical study of training end-to-end vision-and-language
  transformers.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition}, pp.\  18166--18176, 2022.

\bibitem[Fu et~al.(2021)Fu, Li, Gan, Lin, Wang, Wang, and Liu]{fu2021violet}
Tsu-Jui Fu, Linjie Li, Zhe Gan, Kevin Lin, William~Yang Wang, Lijuan Wang, and
  Zicheng Liu.
\newblock Violet: End-to-end video-language transformers with masked
  visual-token modeling.
\newblock 2021.

\bibitem[Huang et~al.(2019)Huang, Le~Bras, Bhagavatula, and
  Choi]{huang2019cosmos}
Lifu Huang, Ronan Le~Bras, Chandra Bhagavatula, and Yejin Choi.
\newblock Cosmos qa: Machine reading comprehension with contextual commonsense
  reasoning.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pp.\  2391--2401, 2019.

\bibitem[Iki \& Aizawa(2021)Iki and Aizawa]{iki2021effect}
Taichi Iki and Akiko Aizawa.
\newblock Effect of visual extensions on natural language understanding in
  vision-and-language models.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pp.\  2189--2196, 2021.

\bibitem[Kenton \& Toutanova(2019)Kenton and Toutanova]{kenton2019bert}
Jacob Devlin Ming-Wei~Chang Kenton and Lee~Kristina Toutanova.
\newblock Bert: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock In \emph{Proceedings of NAACL-HLT}, pp.\  4171--4186, 2019.

\bibitem[Krishna et~al.(2017)Krishna, Zhu, Groth, Johnson, Hata, Kravitz, Chen,
  Kalantidis, Li, Shamma, et~al.]{krishna2017visual}
Ranjay Krishna, Yuke Zhu, Oliver Groth, Justin Johnson, Kenji Hata, Joshua
  Kravitz, Stephanie Chen, Yannis Kalantidis, Li-Jia Li, David~A Shamma, et~al.
\newblock Visual genome: Connecting language and vision using crowdsourced
  dense image annotations.
\newblock \emph{International journal of computer vision}, 123:\penalty0
  32--73, 2017.

\bibitem[Li et~al.(2022{\natexlab{a}})Li, Li, Li, Niebles, and
  Hoi]{li2022align}
Dongxu Li, Junnan Li, Hongdong Li, Juan~Carlos Niebles, and Steven~CH Hoi.
\newblock Align and prompt: Video-and-language pre-training with entity
  prompts.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition}, pp.\  4953--4963, 2022{\natexlab{a}}.

\bibitem[Li et~al.(2021)Li, Selvaraju, Gotmare, Joty, Xiong, and
  Hoi]{li2021align}
Junnan Li, Ramprasaath Selvaraju, Akhilesh Gotmare, Shafiq Joty, Caiming Xiong,
  and Steven Chu~Hong Hoi.
\newblock Align before fuse: Vision and language representation learning with
  momentum distillation.
\newblock \emph{Advances in neural information processing systems},
  34:\penalty0 9694--9705, 2021.

\bibitem[Li et~al.(2022{\natexlab{b}})Li, Li, Xiong, and Hoi]{li2022blip}
Junnan Li, Dongxu Li, Caiming Xiong, and Steven~CH Hoi.
\newblock Blip: Bootstrapping language-image pre-training for unified
  vision-language understanding and generation.
\newblock 2022{\natexlab{b}}.

\bibitem[Li et~al.(2019)Li, Yatskar, Yin, Hsieh, and Chang]{li2019visualbert}
Liunian~Harold Li, Mark Yatskar, Da~Yin, Cho-Jui Hsieh, and Kai-Wei Chang.
\newblock Visualbert: Asimple and performant baseline for vision and language.
\newblock \emph{arXiv preprint arXiv:1908.03557}, 2019.

\bibitem[Li et~al.(2020)Li, Yin, Li, Zhang, Hu, Zhang, Wang, Hu, Dong, Wei,
  et~al.]{li2020oscar}
Xiujun Li, Xi~Yin, Chunyuan Li, Pengchuan Zhang, Xiaowei Hu, Lei Zhang, Lijuan
  Wang, Houdong Hu, Li~Dong, Furu Wei, et~al.
\newblock Oscar: Object-semantics aligned pre-training for vision-language
  tasks.
\newblock In \emph{Computer Vision--ECCV 2020: 16th European Conference,
  Glasgow, UK, August 23--28, 2020, Proceedings, Part XXX 16}, pp.\  121--137.
  Springer, 2020.

\bibitem[Lin et~al.(2014)Lin, Maire, Belongie, Hays, Perona, Ramanan,
  Doll{\'a}r, and Zitnick]{lin2014microsoft}
Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva
  Ramanan, Piotr Doll{\'a}r, and C~Lawrence Zitnick.
\newblock Microsoft coco: Common objects in context.
\newblock In \emph{Computer Vision--ECCV 2014: 13th European Conference,
  Zurich, Switzerland, September 6-12, 2014, Proceedings, Part V 13}, pp.\
  740--755. Springer, 2014.

\bibitem[Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis,
  Zettlemoyer, and Stoyanov]{liu2019roberta}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer
  Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov.
\newblock Roberta: A robustly optimized bert pretraining approach.
\newblock \emph{arXiv preprint arXiv:1907.11692}, 2019.

\bibitem[Madasu \& Rao(2019)Madasu and Rao]{madasu2019sequential}
Avinash Madasu and Vijjini~Anvesh Rao.
\newblock Sequential learning of convolutional features for effective text
  classification.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pp.\  5658--5667, 2019.

\bibitem[Madasu et~al.(2022)Madasu, Oliva, and Bertasius]{madasu2022learning}
Avinash Madasu, Junier Oliva, and Gedas Bertasius.
\newblock Learning to retrieve videos by asking questions.
\newblock In \emph{Proceedings of the 30th ACM International Conference on
  Multimedia}, pp.\  356--365, 2022.

\bibitem[Madasu et~al.(2023)Madasu, Aflalo, Ben Melech~Stan, Tseng, Bertasius,
  and Lal]{madasu2023improving}
Avinash Madasu, Estelle Aflalo, Gabriela Ben Melech~Stan, Shao-Yen Tseng, Gedas
  Bertasius, and Vasudev Lal.
\newblock Improving video retrieval using multilingual knowledge transfer.
\newblock In \emph{Advances in Information Retrieval: 45th European Conference
  on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2--6, 2023,
  Proceedings, Part I}, pp.\  669--684. Springer, 2023.

\bibitem[Mueller \& Thyagarajan(2016)Mueller and
  Thyagarajan]{mueller2016siamese}
Jonas Mueller and Aditya Thyagarajan.
\newblock Siamese recurrent architectures for learning sentence similarity.
\newblock In \emph{Proceedings of the AAAI conference on artificial
  intelligence}, volume~30, 2016.

\bibitem[Ordonez et~al.(2011)Ordonez, Kulkarni, and Berg]{ordonez2011im2text}
Vicente Ordonez, Girish Kulkarni, and Tamara Berg.
\newblock Im2text: Describing images using 1 million captioned photographs.
\newblock \emph{Advances in neural information processing systems}, 24, 2011.

\bibitem[Radford et~al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal,
  Sastry, Askell, Mishkin, Clark, et~al.]{radford2021learning}
Alec Radford, Jong~Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh,
  Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark,
  et~al.
\newblock Learning transferable visual models from natural language
  supervision.
\newblock In \emph{International Conference on Machine Learning}, pp.\
  8748--8763. PMLR, 2021.

\bibitem[Sakaguchi et~al.(2021)Sakaguchi, Bras, Bhagavatula, and
  Choi]{sakaguchi2021winogrande}
Keisuke Sakaguchi, Ronan~Le Bras, Chandra Bhagavatula, and Yejin Choi.
\newblock Winogrande: An adversarial winograd schema challenge at scale.
\newblock \emph{Communications of the ACM}, 64\penalty0 (9):\penalty0 99--106,
  2021.

\bibitem[Sanh et~al.(2019)Sanh, Debut, Chaumond, and Wolf]{sanh2019distilbert}
Victor Sanh, Lysandre Debut, Julien Chaumond, and Thomas Wolf.
\newblock Distilbert, a distilled version of bert: smaller, faster, cheaper and
  lighter.
\newblock \emph{arXiv preprint arXiv:1910.01108}, 2019.

\bibitem[Sap et~al.(2019)Sap, Rashkin, Chen, Le~Bras, and Choi]{sap2019social}
Maarten Sap, Hannah Rashkin, Derek Chen, Ronan Le~Bras, and Yejin Choi.
\newblock Social iqa: Commonsense reasoning about social interactions.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pp.\  4463--4473, 2019.

\bibitem[Sharma et~al.(2018)Sharma, Ding, Goodman, and
  Soricut]{sharma2018conceptual}
Piyush Sharma, Nan Ding, Sebastian Goodman, and Radu Soricut.
\newblock Conceptual captions: A cleaned, hypernymed, image alt-text dataset
  for automatic image captioning.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pp.\  2556--2565,
  2018.

\bibitem[Singh et~al.(2022)Singh, Hu, Goswami, Couairon, Galuba, Rohrbach, and
  Kiela]{singh2022flava}
Amanpreet Singh, Ronghang Hu, Vedanuj Goswami, Guillaume Couairon, Wojciech
  Galuba, Marcus Rohrbach, and Douwe Kiela.
\newblock Flava: A foundational language and vision alignment model.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition}, pp.\  15638--15650, 2022.

\bibitem[Wang et~al.(2018)Wang, Singh, Michael, Hill, Levy, and
  Bowman]{wang2018glue}
Alex Wang, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel~R
  Bowman.
\newblock Glue: A multi-task benchmark and analysis platform for natural
  language understanding.
\newblock In \emph{International Conference on Learning Representations}, 2018.

\bibitem[Wang et~al.(2019)Wang, Pruksachatkun, Nangia, Singh, Michael, Hill,
  Levy, and Bowman]{wang2019superglue}
Alex Wang, Yada Pruksachatkun, Nikita Nangia, Amanpreet Singh, Julian Michael,
  Felix Hill, Omer Levy, and Samuel Bowman.
\newblock Superglue: A stickier benchmark for general-purpose language
  understanding systems.
\newblock \emph{Advances in neural information processing systems}, 32, 2019.

\bibitem[Wang \& Manning(2012)Wang and Manning]{wang2012baselines}
Sida~I Wang and Christopher~D Manning.
\newblock Baselines and bigrams: Simple, good sentiment and topic
  classification.
\newblock In \emph{Proceedings of the 50th Annual Meeting of the Association
  for Computational Linguistics (Volume 2: Short Papers)}, pp.\  90--94, 2012.

\bibitem[Wang et~al.(2021)Wang, Yu, Yu, Dai, Tsvetkov, and Cao]{wang2021simvlm}
Zirui Wang, Jiahui Yu, Adams~Wei Yu, Zihang Dai, Yulia Tsvetkov, and Yuan Cao.
\newblock Simvlm: Simple visual language model pretraining with weak
  supervision.
\newblock In \emph{International Conference on Learning Representations}, 2021.

\bibitem[Williams et~al.(2018)Williams, Nangia, and Bowman]{williams2018broad}
Adina Williams, Nikita Nangia, and Samuel Bowman.
\newblock A broad-coverage challenge corpus for sentence understanding through
  inference.
\newblock In \emph{Proceedings of the 2018 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pp.\  1112--1122, 2018.

\bibitem[Zellers et~al.(2019)Zellers, Holtzman, Bisk, Farhadi, and
  Choi]{zellers2019hellaswag}
Rowan Zellers, Ari Holtzman, Yonatan Bisk, Ali Farhadi, and Yejin Choi.
\newblock Hellaswag: Can a machine really finish your sentence?
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pp.\  4791--4800, 2019.

\bibitem[Zellers et~al.(2021)Zellers, Lu, Hessel, Yu, Park, Cao, Farhadi, and
  Choi]{zellers2021merlot}
Rowan Zellers, Ximing Lu, Jack Hessel, Youngjae Yu, Jae~Sung Park, Jize Cao,
  Ali Farhadi, and Yejin Choi.
\newblock Merlot: Multimodal neural script knowledge models.
\newblock \emph{Advances in Neural Information Processing Systems},
  34:\penalty0 23634--23651, 2021.

\end{thebibliography}

\documentclass{article} % For LaTeX2e
\usepackage{iclr2023_conference,times}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}
\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage[utf8]{inputenc}
\usepackage{hyperref}
\usepackage{url}
\usepackage{multirow}
\usepackage{multicol}
\hypersetup{
    colorlinks=true,
    linkcolor=cyan
    }
\title{Is multimodal vision supervision beneficial to language?}

\author{Avinash Madasu  \\
Department of Computer Science\\
UNC Chapel Hill, USA \\
\texttt{avinashmadasu17@gmail.com} \\
\And
Vasudev Lal \\
Cognitive Computing Research \\
Intel Labs, 
USA\\
\texttt{vasudev.lal@intel.com}
}

\iclrfinalcopy % Uncomment for camera-ready version, but NOT for submission.
\begin{document}

\maketitle

\begin{abstract}
Vision (image and video) - Language (VL) pre-training is the recent popular paradigm that achieved state-of-the-art results on multi-modal tasks like image-retrieval, video-retrieval, visual question answering etc. These models are trained in an unsupervised way and greatly benefit from the complementary modality supervision. In this paper, we explore if the language representations trained using vision supervision perform better than vanilla language representations on Natural Language Understanding and commonsense reasoning benchmarks. We experiment with a diverse set of image-text models such as ALBEF, BLIP, METER and video-text models like ALPRO, Frozen-in-Time (FiT), VIOLET.  We compare the performance of language representations of stand-alone text encoders of these models to the language representations of text encoders learnt through vision supervision. Our experiments suggest that vanilla language representations show superior performance on most of the tasks. These results shed light on the current drawbacks of the vision-language models. The code is available at \url{https://github.com/avinashsai/MML}
\end{abstract}

\section{Introduction}
Vision-language (VL) pre-training ~\cite{radford2021learning,li2021align,li2022blip,bain2021frozen,fu2021violet} has shown tremendous success in the areas of image-text retrieval ~\cite{li2021align, li2022blip}, visual question answering ~\cite{wang2021simvlm,dou2022empirical}, video retrieval ~\cite{bain2021frozen, fu2021violet, madasu2022learning, madasu2023improving}. These models benefit from the mutual supervision of vision and language leading to the superior results on multi-modal tasks. So, the natural question arises: \textit{``Are vision supervised language representations beneficial compared to vanilla language representations on Natural Language Understanding (NLU) tasks?"} 
To understand this, we conduct a study comparing the language representations trained using only the text to the language representations trained using vision supervision. More specifically, we compare the performance of the text encoders used in vision-language models to the vanilla pre-trained text encoders. 

Few works ~\cite{iki2021effect,singh2022flava} evaluated the performance of vision-language and vanilla language models on GLUE. However, there exists a data discrepancy as these models are pre-trained on different domains of data making the comparisons unfair. To overcome this, we pre-train all the vanilla language models with the text captions used in multi-modal pre-training while keeping the identical training setting. Therefore, the only difference in training between vision-language and vanilla language models is the use of vision data.

For our experiments we use a diverse set of image-text models: ALBEF ~\cite{li2021align}, BLIP ~\cite{li2022blip} and METER ~\cite{dou2022empirical} and video-text models: ALPRO ~\cite{li2022align}, Frozen-in-time (FiT) ~\cite{bain2021frozen} and VIOLET ~\cite{fu2021violet}. We evaluate these models on NLU benchmarks GLUE ~\cite{wang2018glue}, Superglue ~\cite{wang2019superglue} and Common sense reasoning datasets such as SocialIQA ~\cite{sap2019social}, CosmosQA ~\cite{huang2019cosmos}, WinoGrande ~\cite{sakaguchi2021winogrande}, CODAH ~\cite{chen2019codah} and HellaSwag ~\cite{zellers2019hellaswag}. 

Our experiments show that (i) vision supervised language representations under perform compared to vanilla language representations on most of the Natural Language Understanding tasks like Natural Language Inference (NLI), sentence similarity, reading comprehension, linguistic probe and textual entailment. (ii) A similar trend is observed for commonsense reasoning benchmarks.
\section{Related Work}
Over the recent years there has been a tremendous progress in training vision and language together using large-scale multi-modal data. ~\cite{li2019visualbert,chenuniter,li2020oscar}. These models combine both the modalities into a single input and are trained using objectives similar to masked language modelling. Another line of work ~\cite{radford2021learning, li2021align, li2022blip,bain2021frozen} explore dual stream architectures in which there is a separate encoder for each of the modalities and the final representations are minimized using contrastive loss.

Natural Language Understanding involves several tasks such as text classification ~\cite{wang2012baselines, madasu2019sequential}, sentence similarity ~\cite{mueller2016siamese}, Natural Language Inference ~\cite{williams2018broad} etc. However to evaluate the capability of models towards a broad range of NLU tasks, benchmarks such as GLUE ~\cite{wang2018glue}, Superglue ~\cite{wang2019superglue} are introduced. Since then, these benchmarks are being used to comprehensively evaluate the performance of language models.
\begin{table}
\caption{Comparison among different image-text and video-text models in-terms of pre-training data, architecture of the text encoders and size of the text encoder. CC denotes Conceptual captions ~\cite{bain2021frozen}, SBU denotes SBU captions ~\cite{ordonez2011im2text} and VG represents visual genome ~\cite{krishna2017visual}.}
\label{models-pretrain-data}
\begin{center}
\begin{tabular}{|l|l|l|l|l|}
\hline
\bf Type  &\bf Model
&\bf Pre-training Data 
&\bf Text Encoder & \bf Num. layers
\\ \hline \\
\multirow{3}{*}{Image-text} & ALBEF  & CC12M + COCO + SBU + VG (14M) & BERT  & 6 \\
& BLIP  & CC12M + COCO + SBU + VG (14M) & BERT  & 12 \\

& METER & CC3M + SBU + VG (4M) & RoBERTa & 6 \\
\hline
\multirow{3}{*}{Video-text} & ALPRO & CC3M + WebVid-2M (5M) & BERT & 6 \\
& FiT & CC3M + WebVid-2M (5M) &  DistilBERT & 6 \\
& VIOLET & YT180M + CC3M + WebVid-2M (11M) & BERT & 12 \\
\hline
\end{tabular}
\end{center}
\end{table}

\section{Experiments}
\subsection{Models}
We experiment with a diverse set of image-text and video-text models. These models differ in the type of pre-training data used, in the architecture of the text encoder and in the sizes the text encoder. The comparison among the models is shown in the table ~\ref{models-pretrain-data}.
\subsubsection{ALBEF}
ALBEF ~\cite{li2021align} is an image-text model pretrained on conceptual captions 12M (CC12M) ~\cite{sharma2018conceptual}, COCO ~\cite{lin2014microsoft}, SBU captions ~\cite{ordonez2011im2text} and visual genome ~\cite{krishna2017visual}. It's text encoder has a pre-trained BERT ~\cite{kenton2019bert} architecture with six transformer encoder layers.
\subsubsection{BLIP}
BLIP ~\cite{li2022align} is proposed as an extension to ALBEF model pretrained using the same data albeit with a large text encoder. It's text encoder has the same configuration as pre-trained BERT.
\subsubsection{METER}
METER ~\cite{dou2022empirical} is an image-text model pretrained on conceptual captions 3M (CC3M),  SBU captions and visual genome. Pre-trained RoBERTa ~\cite{liu2019roberta} with six transformer encoder layers is used as the text encoder. 
\subsubsection{ALPRO}
ALPRO ~\cite{li2021align} is a video-text model whose text encoder has a pre-trained BERT architecture with six transformer encoders. It is pre-trained on a combined data of conceptual captions 3M (CC3M) and WebVid-2M ~\cite{bain2021frozen}.
\subsubsection{Frozen-in-time (FiT)}
Frozen-in-time ~\cite{bain2021frozen} is a dual stream transformer model pre-trained on both image data  conceptual captions 3M (CC3M) and video data WebVid-2M. DistillBERT ~\cite{sanh2019distilbert} is used as the text encoder.
\subsubsection{VIOLET}
VIOLET ~\cite{fu2021violet} is a multi-modal transformer model pre-trained end-to-end on YouTube 180M (YT180M) ~\cite{zellers2021merlot}, conceptual captions 3M (CC3M) and WebVid-2M. The text encoder follows the BERT architecture.
\subsection{Datasets}
For our analysis, we use GLUE, Superglue and commonsense reasoning datasets such as SocialQA, CosmosQA, WinoGrande, CODAH and HellaSwag. For all these datasets, we evaluate the models on the dev data.
\subsection{Implementation}
For fair comparison between the vision supervised text models and vanilla text models, we pre-train the vanilla text models with the text captions from the datasets used for large scale training of image-text and video-text models. Now, the only difference between these models is the use of vision data. We pre-train vanilla text models in the exact setup as vision-language models. We then fine-tune both the vision supervised text models and vanilla text models on downstream tasks. For GLUE, the maximum sentence length used is 200 and the models are trained for 5 epochs. In case of superglue, 250 is the maximum sentence length and the model are trained for 25 epochs. For commonsense reasoning, the models are trained for 10 epochs and 300 is the maximum sentence length. Unless otherwise stated, the results reported are the average of 5 runs.

\begin{table}[!h]
\caption{Results on GLUE benchmark. MNLI-mis refers to the task MNLI mismatched and WNLI denotes the Winograd Schema Challenge. We see that language representations learnt through vision supervision under performs compared to vanilla language representations on all the tasks except WNLI.}
\label{glue}
\begin{center}
\begin{tabular}{|l|l|l|l|l|l|l|l|l|l|}
\hline
  \bf Model &\bf Type &\bf
MNLI &\bf MNLI-mis	&\bf QQP	&\bf SST2	&\bf MRPC &\bf	CoLA &\bf	RTE &\bf	WNLI \\
\hline
\multirow{2}{*}{ALBEF} & Text & 82.77 & 82.68 & 90.54 & 91.44 & 72.81 & 81.50 & 58.12 & 46.01 \\
& Image-text & 61.38 & 61.68 & 79.02 & 80.39 & 66.49 & 69.13 & 50.30 & 56.34 \\
\hline
\multirow{2}{*}{BLIP} & Text & 83.04 & 82.70 & 90.54 & 91.44 & 72.81 & 81.50 & 58.12 & 46.01 \\
 & Image-text & 61.38 & 61.68 & 79.02 & 80.39 & 66.49 & 69.13 & 50.30 & 56.34 \\
\hline
\multirow{2}{*}{METER} & Text & 86.59 & 86.15 & 90.99 & 93.27 & 76.06 & 82.58 & 64.02 & 56.34 \\
 & Image-text & 31.82 & 31.82 & 77.91 & 81.12 & 66.49 & 69.13 & 47.29 & 56.34 \\
\hline
\hline
\multirow{2}{*}{ALPRO} & Text & 82.96 & 82.81 & 90.64 & 92.05 & 70.96 & 79.93 & 60.41 & 45.07 \\
 & Video-text & 62.53 & 63.26 & 79.35 & 80.96 & 66.49 & 69.13 & 54.39 & 56.34 \\
 \hline
 \multirow{2}{*}{FiT} & Text & 79.10 & 80.23 & 89.51 & 52.03 & 72.58 & 69.13 & 57.28 & 48.83 \\
 & Video-text & 59.54 & 59.45 & 79.01 & 52.18 & 66.78 & 69.13 & 48.01 & 56.34 \\
 \hline
 \multirow{2}{*}{VIOLET} & Text & 83.19 & 83.59 & 90.68 & 92.74 & 71.92 & 81.66 & 59.93 & 52.58 \\
 & Video-text & 61.38 & 61.68 & 79.02 & 80.39 & 66.49 & 69.13 & 50.30 & 56.34 \\
\hline
\end{tabular}
\end{center}
\end{table}

\begin{table}[!h]
\caption{Results on Superglue benchmark. WiC represents Word-in-Context, CB represents CommitmentBank, COPA denotes Choice of Plausible Alternatives and WSC means The Winograd Schema Challenge.}
\label{superglue}
\begin{center}
\begin{tabular}{|l|l|l|l|l|l|l|}
\hline
\bf Model &	\bf Type &	\bf BoolQ & \bf	WiC	& \bf CB	& \bf COPA & \bf WSC \\
\hline
\multirow{2}{*}{ALBEF} & Text & 70.41 & 63.13 & 76.79 & 48.00 & 63.46 \\
 & Image-text & 63.30 & 55.02 & 63.93 & 51.60 & 63.46 \\
\hline
\multirow{2}{*}{BLIP} & Text & 70.41 & 63.13 & 76.43 & 48.00 & 63.46 \\
 & Image-text & 63.30 & 55.02 & 63.93 & 51.60 & 63.46 \\
\hline
\multirow{2}{*}{METER} & Text & 72.40 & 66.11 & 75.00 & 46.80 & 63.46 \\
 & Image-text & 66.87 & 53.98 & 69.64 & 50.80 & 63.46 \\
\hline
\hline
\multirow{2}{*}{ALPRO} & Text & 71.16 & 67.18 & 76.79 & 42.20 & 63.46 \\
 & Video-text & 65.17 & 53.17 & 62.50 & 50.60 & 62.50 \\
\hline
\multirow{2}{*}{FiT} & Text & 68.91 & 62.38 & 69.29 & 44.80 & 63.46 \\
 & Video-text & 64.69 & 53.20 & 70.71 & 53.80 & 63.46 \\
\hline
\multirow{2}{*}{VIOLET} & Text & 63.85 & 57.37 & 66.07 & 56.00 & 63.46 \\
 & Video-text & 63.44 & 54.11 & 63.93 & 52.60 & 63.46 \\
\hline

\end{tabular}
\end{center}
\end{table}

\begin{table}[!h]
\caption{Results on Commonsense reasoning tasks.}
\label{commonsense}
\begin{center}
\begin{tabular}{|l|l|l|l|l|l|l|}
\hline
\bf Model & \bf	Type &\bf SocialQA &\bf	CosmosQA & \bf WinoGrande & \bf	CODAH &\bf HellaSwag \\
\hline
\multirow{2}{*}{ALBEF} & Text & 40.50 & 26.45 & 53.12 & 25.72 & 25.04 \\
 & Image-text & 33.47 & 25.24 & 49.57 & 25.72 & 24.48 \\
\hline
\multirow{2}{*}{BLIP} & Text & 52.27 & 25.72 & 56.88 & 26.02 & 25.24 \\
 & Image-text & 33.47 & 25.24 & 49.57 & 25.72 & 24.48 \\
\hline
\multirow{2}{*}{METER} & Text & 58.39 & 31.32 & 59.59 & 24.40 & 25.04 \\
 & Image-text & 33.47 & 25.00 & 49.57 & 25.72 & 24.48 \\
\hline
\hline
\multirow{2}{*}{ALPRO} & Text & 49.90 & 27.45 & 56.56 & 24.10 & 24.89 \\
 & Video-text & 33.96 & 25.70 & 50.28 & 25.72 & 24.48 \\
\hline
\multirow{2}{*}{FiT} & Text & 45.46 & 30.87 & 56.75 & 25.12 & 26.54 \\
 & Video-text & 33.35 & 25.77 & 50.33 & 24.52 & 24.59 \\
\hline
\multirow{2}{*}{VIOLET} & Text & 43.36 & 33.17 & 57.09 & 24.28 & 25.27 \\
 & Video-text & 33.47 & 25.24 & 49.57 & 25.72 & 24.48 \\
\hline
\end{tabular}
\end{center}
\end{table}

\section{Results}
Table ~\ref{glue} shows the results on GLUE benchmark. From the tables, it is evident that vanilla language representations show superior performance compared to vision supervised language representations on most of the tasks across all the models. The drop in performance is significant for NLI tasks like MNLI and MNLI-mismatched (MNLI-mis). A similar trend is observed for sentence similarity (QQP), sentiment classification (SST2), reading comprehension (MRPC), linguistic probe (CoLA) and textual entailment (RTE). However, we see a huge improvement in performance for the Winograd NLI (WNLI) task.

Table ~\ref{superglue} illustrates the results on superglue benchmark. From the table, we observe that vision supervised language representations under perform compared to vanilla language representations. For the tasks question answering (BoolQ), word in context (WiC), discourse (CB) we see a huge drop in performance. However, we see a significant improvement in performance for the casual reasoning (COPA) task. It is worth-noting that the performance is same for both the vanilla and vision supervised language representations on winograd schema challenge (WSC).

Table ~\ref{commonsense} demonstrates the results on commonsense reasoning datasets. As shown in the table, the performance of vanilla language representations surpass vision supervised language representations. There is a notable difference in performance on SocialQA, CosmosQA, WinoGrande and HellaSwag commonsense tasks. However for the CODA dataset, we observe vision supervised language representations outperform vanilla language representations for METER, ALPRO and VIOLET models.

\section{CONCLUSION AND FUTURE DIRECTIONS}
In this paper we comprehensively evaluated if the vision supervised language representations are beneficial to the language. We experimented with three image-text models ALBEF, BLIP, METER and three video-text models ALPRO, FiT, VIOLET on NLU benchmarks GLUE, superglue and commonsense reasoning tasks. Our experiments showed that vanilla language representations significantly outperform vision supervised language representations on most of the tasks. We believe these findings can shed light on the future directions to improve the vision-language pre-training that is beneficial to understanding the language. 
\bibliography{iclr2023_conference}
\bibliographystyle{iclr2023_conference}

\end{document}